1. A rock is thrown at an angle of 45 degrees off the horizontal with an initial velocity of 30 m/s. What is the height of the ball will reach?

2. A ball is thrown horizontally off a 100m cliff. It lands 80m away. At what speed was it thrown?


1
Expert's answer
2021-05-13T14:26:25-0400

1.


"h=\frac{v_{0y}^2}{2g}=\frac{v_0^2\\sin^2\alpha}{2g}=\frac{30^2\\sin^245°}{2\cdot 9.81}=22.9\ (m)"


2.


"s=v_{0x}t=v_0\cdot \sqrt{\frac{2h}{g}}\to"

"v_0=s/\sqrt{\frac{2h}{g}}=80/\sqrt{\frac{2\cdot100}{9.81}}=17.7\ (m/s)"